OpenStudy (anonymous):

transform the following ppolar equation into an equation in rectangluar coordinates: r=2cos theta A. x+y=2 B. (x-1)^2+y^2=1 c. x=2 d. y=2x

OpenStudy (anonymous):

B?

OpenStudy (mertsj):

multiply both sides by r r^2=2rcos theta

OpenStudy (mertsj):

But x^2+y^2=1 and rcos theta = x so: x^2+y^2 = 2x x^2-2x +1 +y^2 = 1 (x-1)^2 + y^2 = 1

OpenStudy (mertsj):

Yep. B